OpenStudy (anonymous):

If a ball is thrown into the air with a velocity of 30m/s, its height in metres after t seconds is given by y=30t-4.9t^2. a) Find the average velocity for the time period beginning when t=2 and lasting 1 sec. b) Find the instantaneous velocity when t=2.

OpenStudy (jhannybean):

To find average velocity, use the second definition of the derivative. \[\frac{f(b)-f(a)}{b-a}\]

OpenStudy (jhannybean):

\[f(t) = 30t-4.9^2\]\[f(2)=~?\]\[f(1)=~?\] Instantaneous velocity is found by taking the limit as \(h\rightarrow 0\)

OpenStudy (zale101):

t=2 and lasting 1 sec I was thinking f(2)=? f(2+1) which is f(3)=?

OpenStudy (zale101):

And \( Average~ velocity=\Large\frac{f(3)-f(2)}{3-2}\)

OpenStudy (anonymous):

How did you know to add the one?

OpenStudy (zale101):

Because when it said t=2 and lasting 1 sec, there's time after t=2 lasted a second later

OpenStudy (jhannybean):

Imagine the ball going up in 2 seconds, but coming down 3 seconds later.

OpenStudy (jhannybean):

factors such as wind resistance, gravitational pull, etc. You know, lots of minimal factors :P

OpenStudy (anonymous):

oohh i see thanks!